LBC/LUCI Science Preparation 2011 05 22 UT

Observers: Olga Kuhn, Andrew Rakich, Doug Miller (Tucson), and Hinz et al (LBTO) OSA: Steve Allanson (LBTO)

* 18:30 RN tests for LUCI. Nominal. luci.20110522.0001-4.

* LUCI scattered light 300s exposure. 5

* 18:45 Rebooted CMU

* 19:15 Put in calibration unit. Manually turn off/on lamps. Take a few arcs with NB7 "LS0.75_450um" (dome is open and it is twilight. 6 (off) 7 (Xe) 8 (off) 9 (Ar).

* 19:20 Spec flat with NB5 (forgot to change in script). Lamp off 10-12, Halo2 13-15.

* 19:30 Cal unit out of beam. Restore to LUCIFER to EndOfNight configuration

Skyflats

* 19:35 Slew to evening flat field position and take test BV flats (created by mkskyflat1.pl). 23655/23655. Both are saturated.
Try SY combination. 23839/23840 28k/8k. scaling off. 23950,23953 21k, ?k
Back to BV combination 024137/024137 49k, 39k
BV flats, set of 5 dithered positions. 38k, 28k for first pair. 32k, 22k (factor 1.5x).

* Blue_SkyFlatScalings.ob to check scalings against V-Bessel Blue. Sky counts too low.

Tests of Temperature Dependence of Collimation Models

* 19:54 Slew to field near Polaris to collimate. First to WT10_401 to range-balance.
Initial 25849, 25838 B, R before collimation.
Collimate with dofpia,/first. Background still high: 21k Blue and 16k Red and no pupils found in Blue image. Blue image has coma which is not helping. Continue dofpia,/first to see if lower background helps. Red is converged already... Still no pupils on Blue. Looks like SX needs Z7 ~ -1500nm.

* Restart dofpia, /first to see if lower background helps with blue. Yes, found 2 pupils. Background is 8000cnts.

* range balance on WT10_401: 31618/31615 b/r Done. 31907/31904 (rotcen_pa180 to obtain these data files: 32107/32104)

* Slew to Polaris field and collimate with dofpia. LBCTemperature/LBCCollimationTemperature_test.ob (15:30:19.84, +69:58:55) Blue has a lot of defocus and no spherical (low, 43deg elevation, seeing est 2"). Second iteration blue z11/z22 fit (<1.5"). Both converged but SX appears to have astigmatism and converged only for seeing est = 1.7".

* 20:30 Start a long, 30-minute integration at this field. A 5-pt dither with 2 filters per channel. Watching for Active Optics corrections (selected low-throughput filters to avoid AO z4 corrections, since this is a test of collimation). No star measured on first pair of AO tech chip images.

Dither starts with non-zero offset. These are recorded correctly in headers, and TELRA, TELDEC header values reflect these initial offsets.

Red side did send a focus correction at 03:43:26.684 of Z4 = -581.0

Blue guide star on image 034734 looks out of focus, 2.1" seeing. Red guide star looks sharp on 34732, 1.5" seeing.

LBTtools.Sandbox.iqplot on 34358 (Blue) gives average FWHM = 6.875 (1.54") and ellip = 0.129.

20:52. Steve is seeing 1.1" on the DIMM. average FWHM on 35110 (blue) is 7.709 (1.72") and on 35107 (red) is 14.96 (3.35") and ellip=0.495

lbcr.20110522.035107.fits All stars are trailed from lower left to upper right. Not typical of red images in sequence.

iqplot on Blue 35824 gives average FWHM = 8.741 (1.95") so focus is falling apart on Blue side. Red has had 3 z4 corrections sent by active optics.

* 21:04 OB completed. Now running dofpia on same field (15h70d). Initial offsets are:
SX: -3977 -555 -240 -244 92 160 81 1.24"
DX: -138 203 -442 -241 -154 -98 -27 0.94"

* 21:12, there is a X-gradient on the SX mirror, from 0.3 on the inside to -0.4 on the outside (cooler side is by the steel, opposite of the way it usually is). It is warming up outside.

* 21:14 After collimating, repeating experiment. But now B AO chip fit a star

04:23:02.097235 B Trackers ActiveOptics P03 +792.

Quitting this OB at 04:42.

* dofpia Seeing estimates are 1.15 B, 0.86 R.
  • Ran 3 superfoc sequences:
    • Blue 1: FWHM = 3.35 for -0.00226 mm for V-BESSEL. Analysis around RC: 0.006mm FWHM = 3.71
    • Red 1: FWHM=3.05 for 0.068mm at r-SLOAN. Analysis around RC: 0.065mm FWHM=3.17
    • Blue Bessel: 0.041mm for 3.26 at B-BESSEL. Analysis around RC: 0.040 FWHM=3.30
    • Analysis around RC: Red F972N20: 0.145 with FWHM=2.24
    • Blue 2: -0.02mm for second set at V-BESSEL. Analysis around RC: -0.013 mm FWHM = 3.20
    • Red 2: 0.047 mm for second set at r-SLOAN. Analysis around RC: 0.043 mm FWHM=2.53

* 21:58 Collimate. So DX is oscillating in Z4, +/- 300 nm. We saw a focus drift of -200 nm from the first Vr superfoc to the second (with no collimation) in between. Conditions are not that stable, although seeing is good. SX mirror still has a strong X gradient.

*22:00 Changing configuration from LBC to LUCIFER/IRTF (via LMIRCAM to do a 30-minute test). Taking a series of 10 LBC biases now.

*23:00 Configuration to LBTI is complete. There was an MCS problem requiring a call to Chris & Tom. And then an M3 problem.

23:43 Sending a MOS mask (NB30) to the FPU. MOS Background went all RED, but this occurred after BLIND/MASK was in STORAGE and NO MASK was in use. Clicked *Initialize and all is OK with NO MASK/NO MASK IN USE. Leave it at this while we continue checkout. (More MOS problems later, which turned into a situation which required the help of Peter. There were two errors: timeout and clamp X loosing steps, which has been noted and will be attended to during summer shutdown. During this troubleshooting I noted that the Mask GUI Config file did not match the wiki, and did correct it. It sounds like the file could explain why NB30 did not move to the FPU without error, however the timeout and clamp X problems seem to be network and hardware related).

LUCIFER Checkout

* 23:52 Steve reset encoders. It was off by 11" in one direction.

Starting "absorb" tests

* Steve sending a preset to BS9162 PA=0, POSITION, TRACK mode.

* Oops - we will point and collimate first. Guider shows 0.55" and DIMM 1.2"

* I. Back to BS9162. PA=20, POS, TRACK.
  • run off_RD_rel => images number 18, 19.
  • Steve "absorbs" RADEC.
  • dice5_RD_abs. (uses absolute in RADEC). 20-24
  • dice5_XY_abs 25-29

* II. Resend TRACK PA=20 POS preset to BS9162
  • run off_XY_rel => 30, 31
  • Steve "absorbs" RADEC (appears to have done nothing - the DETXY offset will not absorb)
  • dice5_RD_abs 32-36
  • dice5_XY_abs 37-41

* III. Resend preset
  • run _off_XY_rel (42, 43) followed by off_RD_rel (44,45)
  • Steve "absorbs" RADEC
  • dice5_RD_abs (46-50)
  • dice5_XY_abs (51-55)

* IV. Resend preset
  • run off_RD_rel (56,57)
  • Steve "absorbs" DETXY (no offset seen to be absorbed.. Steve sees no effect)
  • dice5_XY_abs (58-62)
  • dice5_RD_abs (63-67)

* V. Resend preset
  • run _off_XY_rel (68,69)
  • Steve "absorbs" DETXY (Steve says it absorbed to pointing origin base).
  • dice5_RD_abs (70-74)
  • dice5_XY_abs (75-79)

* VI. Resend preset
  • run off_XY_rel (80,81) and off_RD_rel (82,83)
  • Steve "absorbs" DETXY
  • dice5_XY_abs (84-88)
  • dice5_RD_abs (89-93)

* VII. Resend preset
  • run off_RD_rel (94,95)
  • Steve "absorbs" DETXY and RADEC
  • dice5_RD_abs (96-100)
  • dice5_XY_abs (101-105)

* VIII. Resend preset
  • run _off_XY_rel (106,107)
  • Steve "absorbs" DETXY and RADEC
  • dice5_RD_abs (108-112)
  • dice5_XY_abs (113-117)

* IX Resend preset
  • run off_RD_rel (118,119) followed by off_XY_rel (120,121)
  • Steve "absorbs" DETXY and RADEC
  • dice5_RD_abs (122-126)
  • dice5_XY_abs (127-131)

Check imaging mode and absolute offsets (though this was done by dice5, this script will also guide and collimate)

  • HS1946H.txt while before I could not bring up the PCSGUI to display, now I cannot bring up the GCSGUI
    • Images luci.20110522.0132.fits - stars appear slightly elongated (6.0, 6.9).
    • HS1946H.txt while before I could not bring up the PCSGUI to display, now I cannot bring up the GCSGUI
    • Images luci.20110522.0132.fits - stars appear slightly elongated (6.0, 6.9).
    • MOS goes into "unknown" state when I try to move it to FPU.

* Contacted Walter and later Peter at 09:19 UT. Peter is looking into this. In the meantime, I noticed that the Mask Config GUI does not match what is on the wiki. Will edit once Peter finishes.

* Andrew called Doug and Doug, Andrew & Steve are executing IRTF observations to measure offsets on rfbg. Seeing is very good, 0.5" so these are the type of conditions they have needed to complete rfbg commissioning.

* Edited Mask Config GUI (sorry I did not check this last night or this afternoon! Thought about it, but did not).

* Peter reports there were two MOS errors: One was a timeout error. The other was an error with the clamp X which had lost 500 steps. He says it has been losing steps more recently, and shows progressing degradation. This will be attended to during the summer shutdown.

* 10:52 I am moving mask number 16 to the FPU. And then to number 6 (1"). OK.

Done with IRTC

Test IRTC LUCIFER binocular range balancing

luci 145 and irtc 113330. I forgot to change the sign on both x and ry... so we try again.
run binorb on these and repeat

149 114418
150 114810
151 114850
153 115318 looks like binorb moved star towards center
154 115701 but now after this pair, binorb moved LUCI star away from center
155 120042

* 12:03 Ks flats start at 6000 counts for 2s

* Series of Ks flats to finish up (usual flat position: zenith, Rot=341, vent doors closed)

* 12:11 Steve closing dome

* 12:12 Putting in optic sieve. N3.75 (176) and N1.8 (177)

* Concerned about some elongation on images (taken at H and Ks, where elongation was in similar sense for both). Elevation was low, about 42, but guide star did not seem elongated. Note however that the "absorb" tests were done in imaging mode at higher elevation. Exposures were shorter but unguided and no elongation was obvious.

* Did not do the rotator test.
* Did not yet clear data directory: 16382Mb free now.

-- OlgaKuhn - 22 May 2011
Topic revision: r4 - 24 May 2011, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback